OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable y is eliminated in the first and third equations, then the first and second equations. 4x + y + 3z = 12 5x - 2y - 2z = -11 -x - y - 6z = -9 (4 points) Answer 4x + 3z = 12 x + 6z = 9 3x - 3z = 3 13x + 4z = 13 3x - 3z = 3 x + 6z = 9 4x + 3z = 12 13x + 4z = 13

OpenStudy (anonymous):

did you want to learn how to do it ? or just the answer?

OpenStudy (anonymous):

whatever easier for you .

OpenStudy (anonymous):

im bored so i dont mind explaining.. i'll leave it up to you :)

OpenStudy (anonymous):

umm , well i want you to explain it

OpenStudy (anonymous):

eq1) 4x + y + 3z = 12 eq3) -x - y - 6z = -9 which ever variable you are eliminating(in this case Y) the coefficient should be equal and of opposite sign. (and same degree) Since the Y variable already meets these conditions.. you just have to add these equations together

OpenStudy (anonymous):

Try it and let me know what you get.

OpenStudy (anonymous):

Also, are you okay with why I picked equations 1 and 3?

OpenStudy (anonymous):

im sorrry can i just get the answer . im doing other problems .

OpenStudy (anonymous):

eq1 added to eq3 gives : 3x - 3z = 3 2x(eq1) added to eq2 gives : 13x + 4z = 13